OpenStudy (anonymous):

if log(subscript a) x=7, what is the value of log(subscript a) (a/x^2) ?

zepdrix (zepdrix):

\[\Large\bf\sf \log_ax=7,\qquad\qquad \log_a\frac{a}{x^2}=?\]

zepdrix (zepdrix):

We can apply some rules of logs to help us out!

OpenStudy (anonymous):

I'm all ears

zepdrix (zepdrix):

Here is an important rule to remember:\[\Large \bf\sf \color{orangered}{\log\left(\frac{b}{c}\right)\quad=\quad \log (b)-\log (c)}\]

zepdrix (zepdrix):

So that will allow us to take the first step in rewriting our expression.\[\Large\bf\sf \log_a\frac{a}{x^2}\quad=\quad \log_aa-\log_ax^2\]

zepdrix (zepdrix):

Understand how I applied the rule there? :o

OpenStudy (anonymous):

I think so :) ... I'm taking notes.

OpenStudy (anonymous):

So where do I go from here?

zepdrix (zepdrix):

Whenever the `base` of a log matches the `contents` of the log, the result is 1. If you don't understand why, then it might seem a little weird. You can punch it into your calculator to check. Like if you try \(\Large\bf\sf log_{10}10\) you should get 1. So we can simplify the first term down.

zepdrix (zepdrix):

\[\Large\bf\sf \log_aa-\log_ax^2\quad=\quad 1-\log_ax^2\]

OpenStudy (anonymous):

oh ok

zepdrix (zepdrix):

Another important log rule:\[\Large\bf\sf \color{royalblue}{\log(b^c)\quad=\quad c \log(b)}\]Allows us to bring an exponent down in front of the log as a coefficient.

zepdrix (zepdrix):

Understand how we can apply this to the 2 on x^2?

OpenStudy (anonymous):

Yes 2 log x

zepdrix (zepdrix):

Ok good! \[\Large\bf\sf 1-2\log_ax\]Lemme color it a sec, make something will jump out at you.\[\Large\bf\sf 1-2\color{#DD4747 }{\log_ax}\]That is our expression so far. Here is the initial information they gave us:\[\Large\bf\sf \color{#DD4747}{\log_ax=7}\]

OpenStudy (anonymous):

the coloring really helps!

zepdrix (zepdrix):

Do you see where we can plug in our 7?

OpenStudy (anonymous):

yes so 1-2(7)

zepdrix (zepdrix):

Yesss, good good good.

OpenStudy (anonymous):

-13?

zepdrix (zepdrix):

Yay good job! \c:/

OpenStudy (anonymous):

Thanks:) Wait so that's the final answer?

zepdrix (zepdrix):

This is a pretty tough problem. You have to use 3 different log rules, then use the initial data. It's a bit of a thinker!

zepdrix (zepdrix):

Ya that should be right :o

OpenStudy (anonymous):

Jeez

OpenStudy (anonymous):

Ok because I thought I had to plug it into log a (a/x^2)

OpenStudy (anonymous):

and I didn't square anything...

zepdrix (zepdrix):

They gave a nice simple number for log_a(x). So the idea was: we wanted to twist things around to try and get a log_a(x) out of that messy thing :o

OpenStudy (anonymous):

ok, so I don't have to use the second part of the problem? sorry I'm just confused :)

zepdrix (zepdrix):

Second part of the problem? You mean this?\[\Large\bf\sf \log_a\frac{a}{x^2}=?\]

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

We did use that silly! That's what we `started` with. And widdled it down doing some steps:\[\Large\bf\sf \log_a\frac{a}{x^2}\quad=\quad\log_aa-\log_ax^2\quad=\quad1-2\log_ax\quad=\quad1-2(7)\]

OpenStudy (anonymous):

Oh wow, yes we did.... my bad :)

OpenStudy (anonymous):

Can you help me with one more problem?

zepdrix (zepdrix):

Mm I can try -_- gotta hit the sack soon. So late!

OpenStudy (anonymous):

Explain the algebraic process of how functions can be shown to be inverses. Use the functions below to verify your explanation. f(x) = 3x-1/x+2 and g(x) = 2x+1/3-x

OpenStudy (anonymous):

Ok I REALLY appreciate all your help! You are a life saver!!! And I know I'm tired too!

zepdrix (zepdrix):

\[\Large\bf\sf f(x)=\frac{3x-1}{x+2},\qquad\qquad g(x)=\frac{2x+1}{3-x}\]Let's fiddle around with f(x) and try to show that g(x) is it's inverse. We'll start by letting f(x) = y.\[\Large\bf\sf y=\frac{3x-1}{x+2}\]

OpenStudy (anonymous):

ok

zepdrix (zepdrix):

To find the inverse of this function, we `swap/trade` the x's and y's in the equation.\[\Large\bf\sf x=\frac{3y-1}{y+2}\]

zepdrix (zepdrix):

To find the inverse function, we now need to solve for y.

OpenStudy (anonymous):

ok

zepdrix (zepdrix):

It'll take a few tricky algebra steps to get our y.

zepdrix (zepdrix):

We'll multiply both sides by (y+2)

zepdrix (zepdrix):

\[\Large\bf\sf x(y+2)=3y-1\]

OpenStudy (anonymous):

ok, with u so far

zepdrix (zepdrix):

We'll distribute the x to each term in the brackets.

OpenStudy (anonymous):

xy + 2x

zepdrix (zepdrix):

Ok good.\[\Large\bf\sf xy+2x=3y-1\]

zepdrix (zepdrix):

Now we want to get both of the `y` terms on the same side. So let's subtract 3y from each side.

OpenStudy (anonymous):

ok

zepdrix (zepdrix):

No no no, let's not do that. Let's subtract xy from each side actually... It'll work out a little nicer.

OpenStudy (anonymous):

-3y + xy + 2x= -1?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so what does our equation look like now?

zepdrix (zepdrix):

So I guess that gives us this for now: \[\Large\bf\sf 2x=3y-xy-1\]

OpenStudy (anonymous):

ok

zepdrix (zepdrix):

We'll add 1 to each side,\[\Large\bf\sf 2x+1=3y-xy\]Then we'll want to `factor out` a y from each term on the right.

OpenStudy (anonymous):

ok

zepdrix (zepdrix):

\[\Large\bf\sf 2x+1=y(3-x)\]Understand the factoring ok?

OpenStudy (anonymous):

pooh yes, I was doing in wrong on my paper, but yes i see

OpenStudy (anonymous):

*oooh

zepdrix (zepdrix):

lol XD dat pooh bear

OpenStudy (anonymous):

:)

zepdrix (zepdrix):

So to get y alone, let's divide the junk over to the other side.

OpenStudy (anonymous):

ok so 2x+1/y = 3-x?

zepdrix (zepdrix):

Woops! We don't want to divide the y to the left side silly! We're trying to get y alone. I was referring to the `(3-x)` as the junk connected to y.

OpenStudy (anonymous):

ok, so what is our equation now?

OpenStudy (anonymous):

2x+1/3-x=y

OpenStudy (anonymous):

got it

zepdrix (zepdrix):

\[\Large\bf\sf \frac{2x+1}{3-x}=y\]Yes, good good. Our final step is to rewrite y as \(\Large\bf\sf f^{-1}(x)\). This new y that we've found now represents the inverse function of f, since we swapped the x's and y's

OpenStudy (anonymous):

oooohhhh and this equals the other equation we were given!!! so they are inverses of each other, right?

zepdrix (zepdrix):

Yesss, good! So we've just found out that \(\Large\bf\sf f^{-1}(x)=g(x)\) Which means f and g are inverse functions of one another! :O

OpenStudy (anonymous):

I cannot even began to thank you enough for all your help!!!!!

zepdrix (zepdrix):

lol np :3 go to bed bunny it's late!! +_+ err at least in this part of the world it is. :P

OpenStudy (anonymous):

No its past midnight here! Thx a million!